aprojectile is projected with speed u at an angle 60 with horizantal from foot of the inclined plane. if the projectile hits the inclined plane horizontally,the range on inclined plane is

Grade:Upto college level

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear akash

projectile hit the plane horizontaly ,it means it hit at its hihest point

H= u2sin260 /2g

and R =u2sin120/g

or R/2 = u2sin120/2g

tan A =H/(R/2)   where A is the angel of plane

        =tan60 /2

         =√3/2

so range on incline plane =H/sinA

 

Put the value of H and SinA  you will get the result


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin